Find the value of the point charge with an electric field of 12 N/C and a distance 3.3 m



1
Expert's answer
2022-02-28T10:06:53-0500

The Coulomb's law says

"E=k\frac{q}{r^2}"

Hence, the charge

"q=Er^2/k\\
=12*3.3^2/9*10^9=1.5*10^{-8}\:\rm C"